Anne's age is seven tenths of Bill's age. In 4 yrs Anne's age will be eight eleventh's of Bill's age. How old is Bill now?
Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

Anne's age is seven tenths of Bill's age. In 4 yrs Anne's age will be eight eleventh's of Bill's age. How old is Bill now?
Let A be Annie's age and B be Bill's age. Four years from now their ages will be A+4 and B+4, respectively.
We know from the first statement above that:
A = .7B
From the second sentence we also know that:
A+4 = (8/11)*(B+4)
Substitue .7B for A in the equation above:
.7B = (8/11)* (B+4)
Solve for B above and then use that value for B in A=.7B to get the value for A.


Answer by richwmiller(17219)

let a= Anne's age
let b= Bill's age
a=(7/10)*b -their ages now
a+4=(8/11)*(b+4) -their ages in four years
Substitute (7/10)*b for a in the second equation
(7/10)*(b)+4=(8/11)*(b+4)
multiply both sides by 110 (10*11)
77*b+440=80*(b+4)
77*b+440=80*b+80*4
77*b+440=80*b+320
subtract 320 from both sides
77b+120=80b
subtract 77b from both sides
120=3b
divide both sides by 3
40=b Bill is 40 years old
a=(7*40)/10
a=7*4=28
check
32=(8/11)*(44)
32=8*4=32
ok